# Accuracy and stability of the accelerated multi-direct-forcing immersed boundary method

### v2

@@ Abstract
The multi-direct-forcing immersed boundary method allows for high accuracy of the no-slip condition in moving-particle problems but suffers from numerical instability if simulation parameters are not carefully chosen. This study investigates the numerical accuracy and stability of the accelerated multi-direct-forcing immersed boundary method. An analysis of the discretized equations of body motion in moving boundary problems identifies a critical parameter that solely determines the numerical stability for the body motion. Additionally, numerical simulations reveal the optimal acceleration parameter that minimizes the error in the no-slip condition and is independent of details of the boundary discretisation, the boundary shape, and spatial dimensionality. This study provides a guideline for establishing numerically stable simulations of moving boundary problems at optimal accuracy of the no-slip condition.
@@ 1 Introduction
In a diffuse-interface scheme, the boundary points are treated as Lagrangian points that move along with the rigid body, allowing them to be defined independently of the background grid. The volume force is distributed to the surrounding grid points using a weighting function, typically a regularized delta function, as originally proposed by Peskin [9]. Uhlmann [10] was the first to introduce a direct-forcing IBM based on this concept. Since the algorithm of a diffuse-interface scheme is not tied to the boundary’s shape or whether it is stationary or moving, the diffuse-interface scheme provides a straightforward framework for handling complex moving boundary problems. However, it generally suffers from low accuracy of the no-slip condition due to the required velocity interpolation and force spreading procedures.
@@
To improve the accuracy of the no-slip condition in the diffuse-interface direct-forcing IBM, Wang et al. [11] proposed the multi-direct-forcing IBM which iteratively calculates the volume force to reduce the error of the no-slip condition. This method successfully reduced the error of the no-slip condition by an order of magnitude after ten iterations, but at a higher computational cost. Wu and Shu [12] formulated a matrix equation determining the volume force and proposed the implicit velocity-correction-based IBM which directly solves the matrix equation. This method can reduce the error of the no-slip condition to machine accuracy, but solving the matrix equation greatly increases the computational cost, in particular for moving boundary problems where a large matrix needs to be inverted every time step. Remarkably, the multi-direct-forcing IBM is equivalent to the Richardson iteration for solving the matrix equation in the implicit velocity-correction-based IBM, and its convergence can be accelerated by selecting a suitable acceleration parameter for the Richardson iteration [13, 14] (section 2 details this consideration).
@@
The present study focuses on the accelerated multi-direct-forcing IBM with a single acceleration parameter. Zhang et al. [14] were the first to propose the accelerated multi-direct-forcing IBM with an optimal acceleration parameter and investigated the influences of several factors, such as the shape of the boundary and the type of the weighting function, on the acceleration parameter. Gsell et al. [15, 16] proposed a simple correction of the direct-forcing IBM and successfully reduced the error of the no-slip condition, just by multiplying the volume force with an error factor which depends on the type of the weighting function. It turns out that this error factor is equivalent to the acceleration parameter. The relatively recently introduced accelerated multi-direct-forcing IBM has not been comprehensively investigated. Especially, the influence of the acceleration parameter on the numerical stability has not been clarified. In the present study, we investigate not only the accuracy of the accelerated multi-direct-forcing IBM, but also its numerical stability for moving boundary problems. Importantly, this study presents simple guidelines for establishing stable simulations of moving boundary problems using the accelerated multi-direct-forcing IBM.
@@
The remainder of this paper is structured as follows: Section 2 outlines the accelerated multi-direct-forcing IBM. Section 3 discusses its numerical stability for moving boundary problems. Section 4 summarizes the results for stationary and moving boundary problems and discusses the accuracy and stability of the accelerated multi-direct forcing IBM. Finally, section 5 concludes the study.
@@ 2.1 General remarks on the direct-forcing IBM
where $\bm{U}_{k}$ is the velocity of the boundary point. The objective of the direct-forcing IBM is to determine the volume force $\bm{g}(\bm{x},t+\Delta{t})$ such that it satisfies the no-slip condition (5).
@@ 2.2 Characteristics of the matrix $\mathcal{A}$
In general, the maximum and minimum eigenvalues provide important information of the matrix $\mathcal{A}$, e.g., the spectral radii of $\mathcal{A}$ and $\mathcal{A}^{-1}$. We calculate the maximum and minimum eigenvalues of $\mathcal{A}$ for the circular cylinder. Fig. 2 shows the maximum and minimum eigenvalues of $\mathcal{A}$ for various values of $\Delta{s}$. From this figure, we can see that the maximum eigenvalue $\lambda_{\max}$ is approximately equal to $0.375$ for $\phi_{4}$ and $0.5$ for $\phi_{3}$, when $\Delta{s}$ is smaller than the half width of the support of the weighting function. For $\Delta{s}$ larger than this width, $\lambda_{\max}$ increases with $\Delta{s}$. On the other hand, the minimum eigenvalue $\lambda_{\min}$ is approximately equal to $\lambda_{\max}$ for $\Delta{s}\geq 1$, but it steeply decreases for $\Delta{s}\leq 1$ and tends to 0. When $\lambda_{\min}$ is close to 0, the determinant of $\mathcal{A}$ is also close to 0, making $\mathcal{A}$ ill-conditioned for small $\Delta{s}$.
@@
where the constant $C_{s}$ ($s=4$ or $3$) is given by Eq. (16). The first choice in Eq. (20) corresponds to the conventional multi-direct-forcing IBM [11]. The second choice in Eq. (21) is based on the error factor derived by decomposing the summation of Eq. (9) into the two components which are locally parallel and normal to the boundary [15, 16]. The third choice in Eq. (22) is the optimal relaxation parameter of the Richardson iteration [14].
As mentioned in section 2.2, $\left\|\mathcal{A}\right\|_{\infty}\simeq\lambda_{\max}$, and $\lambda_{\max}\simeq C_{s}$ for $\Delta{s}\leq s/2$. Thus, it is expected that the choices in Eq. (21) and in Eq. (22) result in similar accuracy of the no-slip condition. In the following, we investigate the accuracy and stability of the IBM for various values of the acceleration parameter $\omega$.
@@ 4 Numerical simulations and discussion
In this section, we investigate the accuracy of the no-slip condition through simulations of stationary boundary problems and the stability of the body motion through simulations of moving boundary problems. In the present simulations, we calculate the temporary flow velocity $\bm{u}^{*}$ in Eq. (3) by using the lattice Boltzmann method as shown in A.
@@ 4.1.1 Circular cylinder fixed in a Poiseuille flow
First, we examine the accuracy of the no-slip condition on the surface of a circular cylinder fixed in a Poiseuille flow. The domain size is $L\times H=200\Delta x\times 200\Delta{x}$, and the top and bottom boundaries of the domain are no-slip walls. The left and right boundaries are periodic with an imposed pressure difference $\Delta{p}$ [18], inducing the Poiseuille flow (in the absence of the cylinder) with the mean flow velocity
$U=\frac{H^{2}}{12\rho_{\rm f}\nu}\frac{\Delta{p}}{L}$ (35)
@@
In the present simulations, we set $Re=100$. Other parameters are set as $\rho_{\rm f}=1$, $\nu=0.06~(\Delta{x})^{2}/\Delta{t}$, $\Delta{p}=1.08\times 10^{-4}~(\Delta{x}/\Delta{t})^{2}$, and $U=0.03~\Delta{x}/\Delta{t}$ in simulation units. We run the simulations until the non-dimensional time $t^{*}=tU/H=75$, by which the flow field has reached a steady state and the error of the no-slip condition remains unchanged.
Figure 4: Maximum error of the no-slip condition on the circular cylinder as a function of the iteration count $\ell$ with the weighting function (a) $\phi_{4}$ and (b) $\phi_{3}$ for various values of the acceleration parameter $\omega$.
Fig. 4 shows the maximum error of the no-slip condition on the circular cylinder with $\Delta{s}\simeq 1.0$ ($N=157$) in steady state. For the weighting function $\phi_{4}$, the error decreases monotonically with the number of iterations and significantly depends on the acceleration parameter $\omega$. Notably, the error for $\omega=C_{4}^{-1}$ is about ten times as small as that for $\omega=1.0$ with $\ell=1$, indicating that a good choice of $\omega$ can significantly improve the accuracy of the no-slip condition even without iteration. This trend can also be observed for the weighting function $\phi_{3}$, although the accuracy for $\phi_{3}$ and $\omega=C_{3}^{-1}$ is inferior to that for $\phi_{4}$ and $\omega=C_{4}^{-1}$.
Figure 5: Maximum error of the no-slip condition on the circular cylinder as a function of the acceleration parameter $\omega$ with the weighting function $\phi_{4}$ and $\phi_{3}$ without iteration ($\ell=1$).
In order to find the optimal value of $\omega$, we plot the error without iteration ($\ell=1$) as a function of $\omega$ as shown in Fig. 5. From this figure, we can see that the optimal values of $\omega$ for $\phi_{4}$ and $\phi_{3}$ are approximately equal to $C_{4}^{-1}=8/3$ and $C_{3}^{-1}=2$, respectively. The maximum eigenvalues of the matrix $\mathcal{A}$ for $\phi_{4}$ and $\phi_{3}$ are $\lambda_{\rm max}=0.3787$ and $0.5042$, respectively, and the maximum norms of $\mathcal{A}$ for $\phi_{4}$ and $\phi_{3}$ are $\left\|\mathcal{A}\right\|_{\infty}=0.3857$ and $0.5270$, respectively. Thus, $C_{s}^{-1}\simeq\lambda_{\rm max}^{-1}\simeq\left\|\mathcal{A}\right\|_{\infty}^{-1}$ for both $s=4$ and $s=3$, and $\omega=C_{s}^{-1}$ is almost the optimal choice to minimize the error of the no-slip condition. These findings are consistent with our expectation in section 2.3 that the choices (21) and (22) should result in the same accuracy of the no-slip condition.
@@
Fig. 6 shows the optimal value of $\omega$ as a function of the distance $\Delta{s}$ between the neighboring boundary points. This figure indicates that the optimal value of $\omega$ is approximately equal to $C_{s}^{-1}$, independently of $\Delta{s}$ in the range of $0.4\leq\Delta{s}\leq 1.6$. Notably, the simulations for $\omega\simeq C_{s}^{-1}$ are stable in this range of $\Delta{s}$, which includes the range where $\mathcal{A}$ becomes ill-conditioned, and the maximum errors of the no-slip condition have comparable values for various values of $\Delta{s}$. Therefore, the choice of $\omega=C_{s}^{-1}$ is almost optimal in terms of the accuracy of the no-slip condition independently of $\Delta{s}$.
@@ 4.1.2 Elliptical cylinder fixed in a Poiseuille flow
Next, we examine the accuracy of the no-slip boundary condition on the surface of an elliptical cylinder, whose boundary points are distributed unevenly, fixed in a Poiseuille flow. The computational domain and parameters are the same as those in section 4.1.1, except that an elliptical cylinder (long and short radii are $a=16\Delta{x}$ and $b=8\Delta{x}$, respectively) is located at the center of the domain instead of a circular cylinder. The major axis of the elliptical cylinder is inclined by $\alpha=45^{\circ}$ with respect to the flow axis. The locations of the boundary points of the cylinder are given by
@@
Figure 7: Maximum error of the no-slip condition on an elliptical cylinder with the weighting function $\phi_{4}$: (a) dependence of the iteration count $\ell$ for various values of the acceleration parameter $\omega$; (b) dependence of $\omega$ without iteration ($\ell=1$).
Fig. 7 shows the maximum error of the no-slip condition on the elliptical cylinder in steady state for the weighting function $\phi_{4}$. From Fig. 7(a), we can see the same trend as for the circular cylinder in section 4.1.1, i.e., the error decreases monotonically with the iteration count, and the error for $\omega=C_{4}^{-1}$ is about ten times as small as that for $\omega=1.0$ with $\ell=1$. Also, from Fig. 7(b), we can see that the optimal value of $\omega$ is approximately equal to $C_{4}^{-1}$ in the same way as for the circular cylinder. Therefore, even when the boundary points are unevenly spaced, the choice of $\omega=C_{s}^{-1}$ can optimize the accuracy of the no-slip condition. The weighting function $\phi_{3}$ gives similar results (not shown).

### v3

# Boundary-velocity error and stability of the accelerated multi-direct-forcing immersed boundary method

@@ Abstract
The multi-direct-forcing immersed boundary method allows for a small velocity error of the no-slip condition in moving-particle problems but suffers from numerical instability if simulation parameters are not carefully chosen. This study investigates the boundary-velocity error and numerical stability of the accelerated multi-direct-forcing immersed boundary method. An analysis of the discretized equations of body motion in moving boundary problems identifies a critical parameter that solely determines the numerical stability for the body motion. Additionally, numerical simulations reveal the optimal acceleration parameter that minimizes the velocity error of the no-slip condition and is independent of details of the boundary discretisation, the boundary shape, and spatial dimensionality. This study provides a guideline for establishing numerically stable simulations of moving boundary problems at optimal boundary-velocity error.
@@ 1 Introduction
In a diffuse-interface scheme, the boundary points are treated as Lagrangian points that move along with the rigid body, allowing them to be defined independently of the background grid. The volume force is distributed to the surrounding grid points using a weighting function, typically a regularized delta function, as originally proposed by Peskin [9]. Uhlmann [10] was the first to introduce a direct-forcing IBM based on this concept. Since the algorithm of a diffuse-interface scheme is not tied to the boundary’s shape or whether it is stationary or moving, the diffuse-interface scheme provides a straightforward framework for handling complex moving boundary problems. However, it generally suffers from a large velocity error of the no-slip condition due to the required velocity interpolation and force spreading procedures.
@@
To reduce the velocity error of the no-slip condition in the diffuse-interface direct-forcing IBM, Wang et al. [11] proposed the multi-direct-forcing IBM which iteratively calculates the volume force. This method successfully reduced the boundary-velocity error by an order of magnitude after ten iterations, but at a higher computational cost. Wu and Shu [12] formulated a matrix equation determining the volume force and proposed the implicit velocity-correction-based IBM which directly solves the matrix equation. This method can reduce the boundary-velocity error to machine accuracy, but solving the matrix equation greatly increases the computational cost, in particular for moving boundary problems where a large matrix needs to be inverted every time step. Remarkably, the multi-direct-forcing IBM is equivalent to the Richardson iteration for solving the matrix equation in the implicit velocity-correction-based IBM, and its convergence can be accelerated by selecting a suitable acceleration parameter for the Richardson iteration [13, 14] (section 2 details this consideration).
@@
The present study focuses on the accelerated multi-direct-forcing IBM with a single acceleration parameter. Zhang et al. [14] were the first to propose the accelerated multi-direct-forcing IBM with an optimal acceleration parameter and investigated the influences of several factors, such as the shape of the boundary and the type of the weighting function, on the acceleration parameter. Gsell et al. [15, 16] proposed a simple correction of the direct-forcing IBM and successfully reduced the boundary-velocity error, just by multiplying the volume force with an error factor which depends on the type of the weighting function. It turns out that this error factor is equivalent to the acceleration parameter. The relatively recently introduced accelerated multi-direct-forcing IBM has not been comprehensively investigated. Especially, the influence of the acceleration parameter on the numerical stability has not been clarified. In the present study, we investigate not only the boundary-velocity error of the accelerated multi-direct-forcing IBM, but also its numerical stability for moving boundary problems. Importantly, this study presents simple guidelines for establishing stable simulations of moving boundary problems using the accelerated multi-direct-forcing IBM.
@@
The remainder of this paper is structured as follows: Section 2 outlines the accelerated multi-direct-forcing IBM. Section 3 discusses its numerical stability for moving boundary problems. Section 4 summarizes the results for stationary and moving boundary problems and discusses the boundary-velocity error and stability of the accelerated multi-direct forcing IBM. Finally, section 5 concludes the study.
@@ 2.1 General remarks on the direct-forcing IBM
where $\bm{U}_{k}$ is the velocity of the boundary point. The objective of the direct-forcing IBM is to determine the volume force $\bm{g}(\bm{x},t+\Delta{t})$ such that it satisfies the no-slip condition (5). It is noted that in the sharp-interface direct-forcing IBM such as the ghost-cell method [8], the flow velocity and pressure at lattice points inside the body (namely ghost cells) are directly specified so as to satisfy the no-slip condition. This procedure is equivalent to applying the volume force $\bm{g}(\bm{x},t+\Delta{t})$ only within ghost cells. Thus, the above formulation provides a general description of direct-forcing IBMs, encompassing both sharp-interface and diffuse-interface schemes.
@@ 2.2 Characteristics of the matrix $\mathcal{A}$
In general, the maximum and minimum eigenvalues provide important information of the matrix $\mathcal{A}$, e.g., the spectral radii of $\mathcal{A}$ and $\mathcal{A}^{-1}$. We calculate the maximum and minimum eigenvalues of $\mathcal{A}$ for the circular cylinder. Fig. 2 shows the maximum and minimum eigenvalues of $\mathcal{A}$ for various values of $\Delta{s}$. From this figure, we can see that the maximum eigenvalue $\lambda_{\max}$ is approximately equal to $0.375$ for $\phi_{4}$ and $0.5$ for $\phi_{3}$, when $\Delta{s}$ is smaller than the half width of the support of the weighting function, i.e., $\Delta{s}\leq s/2$. For $\Delta{s}$ larger than this width, $\lambda_{\max}$ increases with $\Delta{s}$. On the other hand, the minimum eigenvalue $\lambda_{\min}$ is approximately equal to $\lambda_{\max}$ for $\Delta{s}\geq 1$, but it steeply decreases for $\Delta{s}\leq 1$ and tends to 0. When $\lambda_{\min}$ is close to 0, the determinant of $\mathcal{A}$ is also close to 0, making $\mathcal{A}$ ill-conditioned for small $\Delta{s}$.
@@
where the constant $C_{s}$ ($s=4$ or $3$) is given by Eq. (16). The first choice in Eq. (20) corresponds to the conventional multi-direct-forcing IBM [11]. The second choice in Eq. (21) is based on the error factor derived by decomposing the summation of Eq. (9) into the two components which are locally parallel and normal to the boundary [15, 16]. The third choice in Eq. (22) is the optimal relaxation parameter of the Richardson iteration [14]. Also, we can find a similarity to the feedback IBM proposed by Goldstein et al. [19] by interpreting $\omega\rho_{\rm f}/\Delta{t}$ as the feedback coefficient $\beta$.
As mentioned in section 2.2, $\left\|\mathcal{A}\right\|_{\infty}\simeq\lambda_{\max}$, and $\lambda_{\max}\simeq C_{s}$ for $\Delta{s}\leq s/2$. Thus, it is expected that the choices in Eq. (21) and in Eq. (22) result in similar boundary-velocity error. In the following, we investigate the boundary-velocity error and stability of the IBM for various values of the acceleration parameter $\omega$.
@@ 4 Numerical simulations and discussion
In this section, we investigate the velocity error of the no-slip condition through simulations of stationary boundary problems and the stability of the body motion through simulations of moving boundary problems. In the present simulations, we calculate the temporary flow velocity $\bm{u}^{*}$ in Eq. (3) by using the lattice Boltzmann method as shown in A. It should be noted that although the density in the LBM exhibits small variations from the constant fluid density $\rho_{\rm f}$ due to weak compressibility, the flow velocity and pressure calculated by the LBM converge to the incompressible Navier-Stokes equations (1) and (2) with the second-order accuracy [17, 18], and the density variation is a negligible error which does not affect the second-order accuracy.
@@ 4.1.1 Circular cylinder fixed in a Poiseuille flow
First, we examine the velocity error of the no-slip condition on the surface of a circular cylinder fixed in a Poiseuille flow. The domain size is $L\times H=200\Delta x\times 200\Delta{x}$, and the top and bottom boundaries of the domain are no-slip walls. The left and right boundaries are periodic with an imposed pressure difference $\Delta{p}$ [18], inducing the Poiseuille flow (in the absence of the cylinder) with the mean flow velocity
$U=\frac{H^{2}}{12\rho_{\rm f}\nu}\frac{\Delta{p}}{L},$ (35)
@@
In the present simulations, we set $Re=100$. Other parameters are set as $\rho_{\rm f}=1$, $\nu=0.06~(\Delta{x})^{2}/\Delta{t}$, $\Delta{p}=1.08\times 10^{-4}~(\Delta{x}/\Delta{t})^{2}$, and $U=0.03~\Delta{x}/\Delta{t}$ in simulation units. We run the simulations until the non-dimensional time $t^{*}=tU/H=75$, by which the flow field has reached a steady state and the boundary-velocity error remains unchanged.
Figure 4: Maximum boundary-velocity error on the circular cylinder as a function of the iteration count $\ell$ with the weighting function (a) $\phi_{4}$ and (b) $\phi_{3}$ for various values of the acceleration parameter $\omega$.
Fig. 4 shows the maximum boundary-velocity error on the circular cylinder with $\Delta{s}\simeq 1.0$ ($N=157$) in steady state. For the weighting function $\phi_{4}$, the error decreases monotonically with the number of iterations and significantly depends on the acceleration parameter $\omega$. Notably, the error for $\omega=C_{4}^{-1}$ is about ten times as small as that for $\omega=1.0$ with $\ell=1$, indicating that a good choice of $\omega$ can significantly reduce the boundary-velocity error even without iteration. This trend can also be observed for the weighting function $\phi_{3}$, although the boundary-velocity error for $\phi_{3}$ and $\omega=C_{3}^{-1}$ is inferior to that for $\phi_{4}$ and $\omega=C_{4}^{-1}$.
Figure 5: Maximum boundary-velocity error on the circular cylinder as a function of the acceleration parameter $\omega$ with the weighting function $\phi_{4}$ and $\phi_{3}$ without iteration ($\ell=1$).
In order to find the optimal value of $\omega$, we plot the error without iteration ($\ell=1$) as a function of $\omega$ as shown in Fig. 5. From this figure, we can see that the optimal values of $\omega$ for $\phi_{4}$ and $\phi_{3}$ are approximately equal to $C_{4}^{-1}=8/3$ and $C_{3}^{-1}=2$, respectively. The maximum eigenvalues of the matrix $\mathcal{A}$ for $\phi_{4}$ and $\phi_{3}$ are $\lambda_{\rm max}=0.3787$ and $0.5042$, respectively, and the maximum norms of $\mathcal{A}$ for $\phi_{4}$ and $\phi_{3}$ are $\left\|\mathcal{A}\right\|_{\infty}=0.3857$ and $0.5270$, respectively. Thus, $C_{s}^{-1}\simeq\lambda_{\rm max}^{-1}\simeq\left\|\mathcal{A}\right\|_{\infty}^{-1}$ for both $s=4$ and $s=3$, and $\omega=C_{s}^{-1}$ is almost the optimal choice to minimize the boundary-velocity error. These findings are consistent with our expectation in section 2.3 that the choices (21) and (22) should result in the same boundary-velocity error.
@@
Fig. 6 shows the optimal value of $\omega$ as a function of the distance $\Delta{s}$ between the neighboring boundary points. This figure indicates that the optimal value of $\omega$ is approximately equal to $C_{s}^{-1}$, independently of $\Delta{s}$ in the range of $0.4\leq\Delta{s}\leq 1.6$. Notably, the simulations for $\omega\simeq C_{s}^{-1}$ are stable in this range of $\Delta{s}$, which includes the range where $\mathcal{A}$ becomes ill-conditioned, and the maximum boundary-velocity errors have comparable values for various values of $\Delta{s}$. Therefore, the choice of $\omega=C_{s}^{-1}$ is almost optimal in terms of the velocity error of the no-slip condition independently of $\Delta{s}$.
Figure 7: Penetration magnitude $Q_{\rm max}$ across the surface of the circular cylinder with $\Delta{s}=1.0$ as a function of the acceleration parameter $\omega$ for the weighting functions $\phi_{4}$ and $\phi_{3}$ without iteration ($\ell=1$); the insets display the streamlines around the circular cylinder.
From another viewpoint, we examine the penetration across the surface of the circular cylinder, which has negative impacts on the flow field around the boundary as well as the force acting on the boundary. In order to evaluate the penetration, we define the penetration magnitude $Q_{\rm max}$ as follows:
$Q_{\rm max}=\frac{\max_{1\leq k\leq N}\{|\psi(\bm{X}_{k})-\psi(\bm{X}_{k+1})|\}}{US/N},$ (37)
where $\psi$ is the stream function defined by $u=\partial\psi/\partial y$ and $v=-\partial\psi/\partial x$. We calculate $\psi$ at lattice points by integrating $u$ and $v$ using the trapezoidal rule and then interpolate it at boundary points $\bm{X}_{k}$ (we set $\bm{X}_{N+1}=\bm{X}_{1}$ here). In general, $|\psi(\bm{X})-\psi(\bm{Y})|$ is equal to the volume flow rate passing between two points $\bm{X}$ and $\bm{Y}$. Thus, the penetration magnitude $Q_{\rm max}$ indicates the maximum volume flow rate passing between neighboring boundary points. Fig. 7 shows the values of $Q_{\rm max}$ as a function of the acceleration parameter $\omega$ for $\Delta{s}=1.0$ without iteration ($\ell=1$). From this figure, we can see that $Q_{\rm max}$ monotonically decreases with increasing $\omega$ both for $\phi_{4}$ and $\phi_{3}$. Notably, unlike the boundary-velocity error shown in Fig. 5, $Q_{\rm max}$ does not have its minimum around the optimal value of $\omega$. In other words, as $\omega$ increases beyond the optimal value, the penetration magnitude becomes smaller while the boundary-velocity error becomes larger. In the streamlines shown in the insets of Fig. 7, penetration occurs across the surface of the circular cylinder for $\omega=1.0$, whereas penetration does not occur for the optimal value of $\omega$ ($\omega=C^{-1}_{4}$ for $\phi_{4}$ and $\omega=C^{-1}_{3}$ for $\phi_{3}$). Therefore, the optimal choice of $\omega$ not only minimizes the boundary-velocity error but also prevents streamlines from penetrating across the boundary even without iteration.
In summary, the acceleration parameter $\omega$ improves the convergence of the iteration, as shown in Fig. 4. Furthermore, there exists an optimal value, $\omega\simeq C_{s}^{-1}$, which minimizes the boundary–velocity error (Fig. 5) and prevents streamlines from penetrating across the boundary (Fig. 7). Remarkably, the present study shows that even without iteration, the optimal choice of $\omega$ yields a boundary-velocity error comparable to that obtained after several iterations in the conventional multi-direct-forcing IBM. Therefore, setting $\omega=C_{s}^{-1}$ and $\ell=1$ is sufficient in most cases, although additional iterations might be used when smaller boundary-velocity errors are desired.
@@ 4.1.2 Elliptical cylinder fixed in a Poiseuille flow
Next, we examine the velocity error of the no-slip boundary condition on the surface of an elliptical cylinder, whose boundary points are distributed unevenly, fixed in a Poiseuille flow. The computational domain and parameters are the same as those in section 4.1.1, except that an elliptical cylinder (long and short radii are $a=16\Delta{x}$ and $b=8\Delta{x}$, respectively) is located at the center of the domain instead of a circular cylinder. The major axis of the elliptical cylinder is inclined by $\alpha=45^{\circ}$ with respect to the flow axis. The locations of the boundary points of the cylinder are given by
@@
Figure 8: Maximum boundary-velocity error on an elliptical cylinder with the weighting function $\phi_{4}$: (a) dependence of the iteration count $\ell$ for various values of the acceleration parameter $\omega$; (b) dependence of $\omega$ without iteration ($\ell=1$).
Fig. 8 shows the maximum boundary-velocity error on the elliptical cylinder in steady state for the weighting function $\phi_{4}$. From Fig. 8(a), we can see the same trend as for the circular cylinder in section 4.1.1, i.e., the error decreases monotonically with the iteration count, and the error for $\omega=C_{4}^{-1}$ is about ten times as small as that for $\omega=1.0$ with $\ell=1$. Also, from Fig. 8(b), we can see that the optimal value of $\omega$ is approximately equal to $C_{4}^{-1}$ in the same way as for the circular cylinder. Therefore, even when the boundary points are unevenly spaced, the choice of $\omega=C_{s}^{-1}$ can optimize the velocity error of the no-slip condition. The weighting function $\phi_{3}$ gives similar results (not shown).